OpenStudy (raquelkelli01):

Find the distance between these points. A (5, 8), B(-3, 4) AB =

OpenStudy (anonymous):

2

OpenStudy (smilie2015):

\[$\begin{gathered} = \sqrt {{{( - 3 - 5)}^2} + {{(4 - 8)}^2}} \hfill \\ = \sqrt {{{( - 8)}^2} + {{( - 4)}^2}} \hfill \\ = \sqrt {64 + 16} \hfill \\ = \sqrt {80} \hfill \\ \end{gathered} $\]

OpenStudy (smilie2015):

using distance formula

OpenStudy (pawanyadav):

AB=[(5-(-3))^2 +(8-4)^2]^1/2 =√(64+16) =√80=4√5